# Analog Foundation Models

### v2

@@ 1 Introduction
Computations performed with AIMC are significantly faster and more energy efficient, but are imprecise due to device-to-device variability, noise, and input/output quantization. Similar to Quantization Aware Training (QAT), methods have been devised to make neural networks more robust to analog noise [27, 28, 29, 30, 31] so that they are amenable for deployment on AIMC accelerators. The efficacy of these methods has been successfully validated on recently developed AIMC chips [19, 20, 21] for small CNNs and RNNs with less than 50 million parameters. However, the robustness of LLMs to analog noise remains uncertain [32], highlighting the need for innovative approaches to improve their resilience without repeating the entire training pipeline. In particular, a key challenge is that, unlike quantization, the noise profile of an AIMC chip is non-deterministic, cannot be efficiently modeled during training, and varies across chips – necessitating a generic training method that can generalize across the diverse noise characteristics of multiple AIMC devices.
@@
Finally, we demonstrate that our models scale better with increased test-time compute compared to quantized models using 4-bit per-channel weight and 8-bit static input quantization, leading to improvements of up to on the MATH-500 benchmark.
@@ 2 Background and Related Work
Three key attributes affect the accuracy of AIMC: weight noise, input DAC quantization and output ADC quantization. Weight noise primarily comes from temporal and device-to-device variations of the NVM conductance that encodes a weight [14] (see figure 1b). Input DAC quantization is equivalent to quantizing the input activations of each layer. However, in contrast to many reduced-precision training papers that assume that the quantization ranges can be dynamically recomputed per-token [36, 37], AIMC typically uses static ranges [30, 38]. Computing the range of a signal dynamically in hardware is very expensive, and therefore will negatively impact the overall efficiency of an AIMC accelerator [39]. Similarly, ADC quantization which quantizes the pre-activation outputs of each layer, must also use static ranges. But although DAC ranges can be configured per-layer, ADC ranges must be identical across layers, as they reflect the fixed dynamic range and precision of the ADCs [38]. Although AIMC hardware advancements can improve these nonidealities, increasing the precision of analog circuits comes at an exponential cost in power and area [40]. Therefore, they cannot be completely eliminated and must be mitigated by adapting the pre-trained neural network with HWA training.
@@ 3.1 Training analog foundation models
We find that although noise injection has been widely believed to provide the most significant improvements in robustness, we demonstrate in further ablation studies (see appendix C) that for LLMs weight clipping yields stronger robustness, and that a combination of the two performs best.
@@ 3.2 Evaluation
To simulate the performance of our models in a more realistic setting, we used a noise model from a state-of-the-art 64-core PCM-based AIMC chip [19]. In this chip, the noise introduced by imprecise programming of the weights into the NVM devices dominates, which is why we chose to focus on this type of noise. As can be seen in figure 8, the amount of noise depends on the conductance state, with higher conductances receiving more noise than lower ones. However, because of an additive noise floor, lower conductance states have a worse signal-to-noise ratio than higher conductance states, with the exception of an exact zero, which we assume to not have any noise. In this paper, we indicate benchmark results obtained with this realistic noise model with . Because the noise model from Le Gallo et al. [19] is specific to one type of NVM device, we also evaluate our foundation models on generic additive Gaussian noise (see eq. 3). Results obtained with this type of noise are indicated with . Unless stated otherwise, experiments involving noise injection during evaluation were repeated 10 times for different random seeds, which we found to be crucial for meaningful comparisons.
@@ 4.1 Analog foundation models are robust to analog noise
When evaluating both models – Phi-3-mini-4k-instruct and Llama-3.2-1B-Instruct – off-the-shelf, we see that only injecting hardware-realistic noise leads to an average drop of and , respectively (see table 1). On more challenging benchmarks such as GSM8K, accuracy drops even more: for Phi-3-mini-4k-instruct and for Llama-3.2-1B-Instruct. Our analog foundation models improve on this significantly, reducing the gap to off-the-shelf FP16 performance to for Phi-3-mini-4k-instruct and for Llama-3.2-1B-Instruct while also including static 8-bit input quantization and globally static 8-bit output quantization. Strong gains can be observed especially for harder tasks such as GSM8K, HellaSwag, and ANLI, where the gaps were reduced by up to . We find that QAT also improves robustness to hardware-realistic analog noise, which is due to the indirect noise injection during training via weight-quantization, and the fact that many weights are rounded to zero, where our model does not apply noise. As table 1 shows, our analog foundation models still outperform models trained with LLM-QAT significantly, especially on harder tasks with differences of up to for Phi-3-mini-4k-instruct and Llama-3.2-1B-Instruct. We also compare our analog foundation models to models that were quantized post-training with SpinQuant [37]. Because the original implementation of SpinQuant uses dynamic per-token input quantization, we report the model performance with it (DI8), along with our implementation using hardware-friendly static input ranges (SI8) which evidently performs worse. In any case, we observe that models quantized using SpinQuant [37] show lower robustness to hardware noise even compared to the original off-the-shelf model.
Model | Benchmarks | Avg. | ||||||||
@@
Analog FM (SI8-W16-O8) | ||||||||||
Analog FM (SI8--O8) | ||||||||||
LLM-QAT (SI8-W4) | ||||||||||
LLM-QAT (SI8-) | ||||||||||
SpinQuant (SI8-W4) | ||||||||||
SpinQuant (SI8-) | ||||||||||
SpinQuant (DI8-W4) | ||||||||||
SpinQuant (DI8-) | ||||||||||
@@
Analog FM (SI8-W16-O8) | ||||||||||
Analog FM (SI8--O8) | ||||||||||
LLM-QAT (SI8–W4) | ||||||||||
LLM-QAT (SI8-) | ||||||||||
SpinQuant (SI8-W4) | ||||||||||
SpinQuant (SI8-) | ||||||||||
SpinQuant (DI8-W4) | ||||||||||
SpinQuant (DI8-) |
@@ 4.3 Deployment of analog foundation models on 4-bit digital hardware
Model | Benchmarks | Avg. | ||||||||
Phi-3-mini-4k-instruct (W16) | ||||||||||
Analog FM+RTN (SI8-W4-O8) | ||||||||||
LLM-QAT (SI8-W4) | ||||||||||
SpinQuant (SI8-W4) | ||||||||||
SpinQuant (DI8-W4) | † | |||||||||
Llama-3.2-1B-Instruct (W16) | ||||||||||
Analog FM+RTN (SI8-W4-O8) | ||||||||||
LLM-QAT (SI8–W4) | ||||||||||
SpinQuant (SI8-W4) | ||||||||||
SpinQuant (DI8-W4) | † | |||||||||
† Models use dynamic per-tensor activation quantization. |
@@ 4.4 Test-time compute scaling
As conventional scaling laws approach their limits, novel directions for scaling LLMs have been explored. In test-time compute scaling, model performance is improved by increasing the amount of compute used by the model at inference time. As AIMC is rather unsuitable for training, but promises orders of magnitudes higher power efficiency for inference, shifting the compute budget from training to inference is an ideal trend for AIMC. Therefore, we investigated whether our models also performed on-par to 4-bit weight quantized models when the compute used for inference is scaled up. To test this, we follow Snell et al. [66] and generate responses to each prompt from the MATH-500 [61] dataset. Using a math process reward model [67, 68], answers are assigned a reward and the best answer is chosen by performing weighted majority voting or by simply picking the answer with the highest reward. For each model, we pick the strategy that performed best. As figure 4 shows, the analog foundation models with noisy weights perform better compared to models trained with 8-bit static input and 4-bit per-channel weight quantization. Additionally, we observe that the gap between the noisy analog foundation model and the original off-the-shelf model decreases as we increase . While this increase is small () for Phi-3-mini-4k-instruct, it is much bigger for Llama-3.2-1B-Instruct (). Interestingly, our results show that the analog foundation models scale better compared to the LLM-QAT models. For Phi-3-mini-4k-instruct, the performance gap between the analog foundation model and the quantized model grows from to . For Llama-3.2-1B-Instruct, the gap grows from to .
@@ 5 Conclusion and Limitations
While our paper lays an important foundation, it also comes with limitations. Training billion-parameter models is resource intensive, especially for larger models. Although the training cost is one-time and our method only requires training on less than 1% of the total pre-training tokens, it might still be infeasible for many researchers to train their own models. However, as our ablation studies on the number of pre-training tokens show, the benefit of increasing the number of tokens diminishes around 20B tokens, and other avenues, such as improving data quality, must be explored to further reduce the performance gap of our models to the FP16 baseline. To reduce training overhead, one could think about employing low-rank adaptations in a HWA manner for training the network, or even better, avoid training altogether and use a post-training method for increasing the robustness of LLMs, similar to PTQ methods. Besides the high resource intensity, another limitation of our work is the relatively large gap in accuracy compared to off-the-shelf models, especially for reasoning tasks such as GSM8K or MATH-500. Further research is needed to reduce this gap. Finally, while we show that our models inherited the safety measures of the original models without significant loss in performance, it should be noted that the risk of LLMs producing toxic or harmful content still persists.
@@ Appendix A Analog RoBERTa
† No noise injection during training was used in the finetuning stage. |
@@ B.1 Data generation
Model | Generation Technique | Benchmarks | Avg. | |||||||
MMLU (5-shot) | BoolQ (0-shot) | Hellaswag (5-shot) | MedQA (2-shot) | AGIEval (0-shot) | Arc-C (10-shot) | Arc-E (10-shot) | ANLI (7-shot) | |||
SI8–W16 | SSS | |||||||||
SI8–W16 | RGS | |||||||||
SI8–W16 | SGS |
@@ B.2 Number of training tokens
Model | Tokens | Benchmarks | Avg. | ||||||||
MMLU (5-shot) | GSM8K (CoT 8-shot) | BoolQ (0-shot) | Hellaswag (5-shot) | MedQA (2-shot) | AGIEval (0-shot) | Arc-C (10-shot) | Arc-E (10-shot) | ANLI (7-shot) | |||
Phi-3-mini-4k-instruct | |||||||||||
Analog FM (SI8–W16–O8) | 1B | ||||||||||
Analog FM (SI8––O8) | 1B | ||||||||||
Analog FM (SI8–W16–O8) | 10B | ||||||||||
Analog FM (SI8––O8) | 10B | ||||||||||
Analog FM (SI8–W16–O8) | 20B | ||||||||||
Analog FM (SI8––O8) | 20B | ||||||||||
Llama-3.2-1B-Instruct | |||||||||||
Analog FM (SI8-W16-O8) | 1B | ||||||||||
Analog FM (SI8–-O8) | 1B | ||||||||||
Analog FM (SI8-W16-O8) | 10B | ||||||||||
Analog FM (SI8–-O8) | 10B | ||||||||||
Analog FM (SI8-W16-O8) | 20B | ||||||||||
Analog FM (SI8–-O8) | 20B | ||||||||||
Analog FM (SI8-W16-O8) | 40B | ||||||||||
Analog FM (SI8–-O8) | 40B |
@@
Model | Tokens | Benchmarks | Avg. | ||||||||
MMLU (5-shot) | GSM8K (CoT 8-shot) | BoolQ (0-shot) | Hellaswag (5-shot) | MedQA (2-shot) | AGIEval (0-shot) | Arc-C (10-shot) | Arc-E (10-shot) | ANLI (7-shot) | |||
LLM-QAT (SI8–W4) | 1B | ||||||||||
LLM-QAT (SI8–W4) | 10B | ||||||||||
LLM-QAT (SI8–W4) | 20B |

### v3

@@ 1 Introduction
Computations performed with AIMC are significantly faster and more energy efficient, but are imprecise due to various device/circuit nonidealities. These include input and output quantization as a result of digital-to-analog and analog-to-digital conversion, as well as device programming noise [27], read noise [28], temperature-dependent conductance variations [29], IR-drop [30, 31], or conductance drift [32, 33]. Similar to Quantization Aware Training (QAT), methods have been devised to make neural networks more robust to analog noise [34, 35, 36, 37, 38] so that they are amenable for deployment on AIMC accelerators. The efficacy of these methods has been successfully validated on recently developed AIMC chips [19, 20, 21] for small CNNs and RNNs with less than 50 million parameters. However, the robustness of LLMs to analog noise remains uncertain [39], highlighting the need for innovative approaches to improve their resilience without repeating the entire training pipeline. In particular, a key challenge is that, unlike quantization, the noise profile of an AIMC chip is non-deterministic, cannot be efficiently modeled during training, and varies across chips – necessitating a generic training method that can generalize across the diverse noise characteristics of multiple AIMC devices.
@@
Finally, we demonstrate that our models benefit equally from increased test-time compute when compared to quantized models using 4-bit per-channel weight and 8-bit static input quantization, even showing improvements of up to on the MATH-500 benchmark.
@@ 2 Background and Related Work
Three key attributes affect the accuracy of AIMC: weight noise, input DAC quantization and output ADC quantization. Weight noise primarily comes from temporal and device-to-device variations of the NVM conductance that encodes a weight [14] (see figure 1b). Static weight noise (does not get resampled during inference) is mainly due to programming noise, which is the conductance error from the target weight that remains after a device has been programmed with an iterative read-write-verify programming scheme [27]. Dynamic weight noise sources whose magnitude vary as a function of time include read noise [28] – due to analog 1/f noise of devices and circuits – and temporal conductance drift, which is the decrease in device conductance over time that is notably observed in PCM devices [33]. Input DAC quantization is equivalent to quantizing the input activations of each layer. However, in contrast to many reduced-precision training papers that assume that the quantization ranges can be dynamically recomputed per-token [43, 44], AIMC typically uses static ranges [37, 45]. Computing the range of a signal dynamically in hardware is very expensive, and therefore will negatively impact the overall efficiency of an AIMC accelerator [46]. Similarly, ADC quantization which quantizes the pre-activation outputs of each layer, must also use static ranges. But although DAC ranges can be configured per-layer, ADC ranges must be identical across layers, as they reflect the fixed dynamic range and precision of the ADCs [45]. Although AIMC hardware advancements can improve these nonidealities, increasing the precision of analog circuits comes at an exponential cost in power and area [47]. Therefore, they cannot be completely eliminated and must be mitigated by on-chip calibration methods and by adapting the pre-trained neural network with HWA training. On-chip calibration methods typically involve calibrating the ADCs to reduce non-linearity and ADC-to-ADC variability, as well as determining per-output scales and offsets to correct for systematic linear nonidealities in each output channel [19]. However, calibration alone is insufficient to achieve high on-chip accuracy because it cannot compensate for the stochastic nature of programming noise, or temporal variations such as read noise and drift. Therefore, adapting the neural network through HWA training is essential for accurate on-chip deployment. Calibration and network adaptation are complementary: calibration reduces deterministic hardware variations, while HWA training makes the network robust to the remaining stochastic and nonlinear effects.
@@ 3.1 Training analog foundation models
We find that although noise injection has been widely believed to provide the most significant improvements in robustness, we demonstrate in further ablation studies (see appendix C) that for LLMs weight clipping yields stronger robustness, and that a combination of the two performs best. This is because clipping applied during training causes smaller weights to be mapped to larger conductance values, which have higher Signal-to-Noise Ratio (SNR) compared to the lower conductance values for the PCM-based model we consider. As a result, the average per-weight SNR increases, resulting in higher robustness (see appendix D for details).
@@ 3.2 Evaluation
To simulate the performance of our models in a more realistic setting, we used a noise model from a state-of-the-art 64-core PCM-based AIMC chip [19]. In this chip, the noise introduced by imprecise programming of the weights into the NVM devices - programming noise - dominates, which is why we chose to focus on this type of noise. Note that programming noise does not originate directly from device-to-device variability, which is mostly accounted for by the iterative read-write-verify programming scheme that iteratively nudges the device conductance towards a target value. Programming noise is actually the conductance error from the target weight that remains after a device has been programmed. As can be seen in figure 9, the amount of noise depends on the conductance state, with higher conductances having more noise than lower ones. However, because of an additive noise floor, lower conductance states have a worse signal-to-noise ratio than higher conductance states. In this paper, we indicate benchmark results obtained with this realistic noise model with . Because the noise model from Le Gallo et al. [19] is specific to one type of NVM device, we also evaluate our foundation models on generic additive Gaussian noise (see eq. 3). Results obtained with this type of noise are indicated with . Unless stated otherwise, experiments involving noise injection during evaluation were repeated 10 times for different random seeds, which we found to be crucial for meaningful comparisons.
@@ 4.1 Analog foundation models are robust to analog noise
When evaluating both models – Phi-3-mini-4k-instruct and Llama-3.2-1B-Instruct – off-the-shelf, we see that only injecting hardware-realistic noise leads to an average drop of and , respectively (see table 1). On more challenging benchmarks such as GSM8K, accuracy drops even more: for Phi-3-mini-4k-instruct and for Llama-3.2-1B-Instruct. Our analog foundation models improve on this significantly, reducing the gap to off-the-shelf FP16 performance to for Phi-3-mini-4k-instruct and for Llama-3.2-1B-Instruct while also including static 8-bit input quantization and globally static 8-bit output quantization (see further results in appendix F where we compare to an analog foundation model trained with 7-bit input quantization). Strong gains can be observed especially for harder tasks such as GSM8K, HellaSwag, and ANLI, where the gaps were reduced by up to . We find that QAT also improves robustness to hardware-realistic analog noise, which is due to the indirect noise injection during training via weight-quantization. As table 1 shows, our analog foundation models still outperform models trained with LLM-QAT significantly, especially on harder tasks with differences of up to for Phi-3-mini-4k-instruct and Llama-3.2-1B-Instruct. We also compare our analog foundation models to models that were quantized post-training with SpinQuant [44]. Because the original implementation of SpinQuant uses dynamic per-token input quantization, we report the model performance with it (DI8), along with our implementation using hardware-friendly static input ranges (SI8) which evidently performs worse. In any case, we observe that models quantized using SpinQuant [44] show lower robustness to hardware noise even compared to the original off-the-shelf model.
| Model | Benchmarks | Avg. | ||||||||||||||||||||||||
@@
| Analog FM (SI8-W16-O8) | ||||||||||||||||||||||||||
| Analog FM (SI8--O8) | ||||||||||||||||||||||||||
| LLM-QAT (SI8-W4) | ||||||||||||||||||||||||||
| LLM-QAT (SI8-) | ||||||||||||||||||||||||||
| SpinQuant (SI8-W4) | ||||||||||||||||||||||||||
| SpinQuant (SI8-) | ||||||||||||||||||||||||||
| SpinQuant (DI8-W4) | ||||||||||||||||||||||||||
| SpinQuant (DI8-) | ||||||||||||||||||||||||||
@@
| Analog FM (SI8-W16-O8) | ||||||||||||||||||||||||||
| Analog FM (SI8--O8) | ||||||||||||||||||||||||||
| LLM-QAT (SI8–W4) | ||||||||||||||||||||||||||
| LLM-QAT (SI8-) | ||||||||||||||||||||||||||
| SpinQuant (SI8-W4) | ||||||||||||||||||||||||||
| SpinQuant (SI8-) | ||||||||||||||||||||||||||
| SpinQuant (DI8-W4) | ||||||||||||||||||||||||||
| SpinQuant (DI8-) | ||||||||||||||||||||||||||
@@
In addition to sweeping generic additive Gaussian noise, we performed further experiments in appendix F, where we evaluate Phi-3-mini-4k-instruct on a noise model extracted from ReRAM devices [21]. Under this overall stronger noise, our analog foundation model shows better performance compared to the LLM-QAT model (see appendix F.1). To better understand the robustness under hardware-realistic noise at different magnitudes, we scaled the PCM-based programming noise by factors of and . Experiments (see appendix F.1) show that our analog foundation model outperforms the runner-up LLM-QAT model by for the scale and for the scale. Finally, to test the robustness to other types of noise, we conducted experiments where we inject read noise and apply conductance drift according to a publicly available noise model based on hardware [73]. Results on Arc-C and MedQA show that the analog foundation model is significantly more robust to drift and read noise compared to both the off-the-shelf model and the model trained with LLM-QAT (see appendix F.2).
Modern AIMC chips have much smaller tiles compared to the weight matrices of modern LLMs. Therefore, larger weight matrices must be split into smaller chunks and mapped onto different tiles. The MVM is then performed by aggregating partial results of the individual tiles using high(er)-precision digital circuitry. Using tiled matrices has a positive impact on accuracy, primarily because weights are now re-scaled per-tile, instead of per-layer, resulting in higher SNR. Experiments on Phi-3-mini-4k-instruct show that off-the-shelf models and models trained with LLM-QAT benefit from this increased SNR, improving average noisy accuracy by 1.76% and 1.11%, respectively. Because the analog foundation models already have tight weight distributions, tiling does not have an effect on the noisy performance (see appendix G for details). Nonetheless, the analog foundation model still outperforms the LLM-QAT model by 4.43%.
@@ 4.3 Deployment of analog foundation models on 4-bit digital hardware
| Model | Benchmarks | Avg. | ||||||||||||||||||||||||
| Phi-3-mini-4k-instruct (W16) | ||||||||||||||||||||||||||
| Analog FM+RTN (SI8-W4-O8) | ||||||||||||||||||||||||||
| LLM-QAT (SI8-W4) | ||||||||||||||||||||||||||
| SpinQuant (SI8-W4) | ||||||||||||||||||||||||||
| SpinQuant (DI8-W4) | † | |||||||||||||||||||||||||
| Llama-3.2-1B-Instruct (W16) | ||||||||||||||||||||||||||
| Analog FM+RTN (SI8-W4-O8) | ||||||||||||||||||||||||||
| LLM-QAT (SI8–W4) | ||||||||||||||||||||||||||
| SpinQuant (SI8-W4) | ||||||||||||||||||||||||||
| SpinQuant (DI8-W4) | † | |||||||||||||||||||||||||
| † Models use dynamic per-tensor activation quantization. | ||||||||||||||||||||||||||
@@ 4.4 Test-time compute scaling
As conventional scaling laws approach their limits, novel directions for scaling LLMs have been explored. In test-time compute scaling, model performance is improved by increasing the amount of compute used by the model at inference time. As AIMC is rather unsuitable for training, but promises orders of magnitudes higher power efficiency for inference, shifting the compute budget from training to inference is an ideal trend for AIMC. Therefore, we investigated whether our models also performed on-par to 4-bit weight quantized models when the compute used for inference is scaled up. To test this, we follow Snell et al. [74] and generate responses to each prompt from the MATH-500 [68] dataset. Using a math process reward model [75, 76], answers are assigned a reward and the best answer is chosen by performing weighted majority voting or by simply picking the answer with the highest reward. For each model, we pick the strategy that performed best. As figure 4 shows, the analog foundation models with noisy weights perform better compared to models trained with 8-bit static input and 4-bit per-channel weight quantization. Additionally, we observe that the gap between the noisy analog foundation model and the original off-the-shelf model decreases as we increase . While this increase is small () for Phi-3-mini-4k-instruct, it is much bigger for Llama-3.2-1B-Instruct ().
@@ 5 Conclusion and Limitations
While our paper lays an important foundation, it also comes with limitations. Training billion-parameter models is resource intensive, especially for larger models. Although the training cost is one-time and our method only requires training on less than 1% of the total pre-training tokens, it might still be infeasible for many researchers to train their own models. Adding to this, input/output quantization and noise injection make the linear kernel approximately three times slower compared to the vanilla version, and, as we show in appendix H, training with noise injection and output quantization leads to slower convergence. As our ablation studies on the number of pre-training tokens show, the benefit of increasing the number of tokens diminishes around 20B tokens, and other avenues, such as improving data quality, must be explored to further reduce the performance gap of our models to the FP16 baseline. To reduce training overhead, one could think about employing low-rank adaptations in a HWA manner for training the network, or even better, avoid training altogether and use a post-training method for increasing the robustness of LLMs, similar to PTQ methods. Besides the high resource intensity, another limitation of our work is the relatively large gap in accuracy compared to off-the-shelf models, especially for reasoning tasks such as GSM8K or MATH-500. Further research is needed to reduce this gap. Finally, while we show that our models inherited the safety measures of the original models without significant loss in performance, it should be noted that the risk of LLMs producing toxic or harmful content still persists.
@@
- 1 Introduction
- 2 Background and Related Work
- 3 Methods
- 4 Results
- 5 Conclusion and Limitations
- A Analog RoBERTa
- B Ablations on the training methodology
- C Ablations on hardware aware training
- D The effect of weight clipping on signal-to-noise ratio
- E Evaluation details
- F Additional noise experiments
- G Impact of tiled AIMC architectures
- H Training convergence
- I Training details
- J Test-time compute scaling details
- K Hardware throughput and energy efficiency estimation
@@ Appendix A Analog RoBERTa
| † No noise injection during training was used in the finetuning stage. | |||||||||
@@ B.1 Data generation
| Model |
Benchmarks | Avg. | ||||||||||||||||||||
| SI8–W16 | SSS | ||||||||||||||||||||||
| SI8–W16 | RGS | ||||||||||||||||||||||
| SI8–W16 | SGS | ||||||||||||||||||||||
@@ B.2 Number of training tokens
| Model | Tokens | Benchmarks | Avg. | |||||||||||||||||||||||
| Phi-3-mini-4k-instruct | ||||||||||||||||||||||||||
| Analog FM (SI8–W16–O8) | 1B | |||||||||||||||||||||||||
| Analog FM (SI8––O8) | 1B | |||||||||||||||||||||||||
| Analog FM (SI8–W16–O8) | 10B | |||||||||||||||||||||||||
| Analog FM (SI8––O8) | 10B | |||||||||||||||||||||||||
| Analog FM (SI8–W16–O8) | 20B | |||||||||||||||||||||||||
| Analog FM (SI8––O8) | 20B | |||||||||||||||||||||||||
| Llama-3.2-1B-Instruct | ||||||||||||||||||||||||||
| Analog FM (SI8-W16-O8) | 1B | |||||||||||||||||||||||||
| Analog FM (SI8–-O8) | 1B | |||||||||||||||||||||||||
| Analog FM (SI8-W16-O8) | 10B | |||||||||||||||||||||||||
| Analog FM (SI8–-O8) | 10B | |||||||||||||||||||||||||
| Analog FM (SI8-W16-O8) | 20B | |||||||||||||||||||||||||
| Analog FM (SI8–-O8) | 20B | |||||||||||||||||||||||||
| Analog FM (SI8-W16-O8) | 40B | |||||||||||||||||||||||||
| Analog FM (SI8–-O8) | 40B | |||||||||||||||||||||||||
@@
| Model | Tokens | Benchmarks | Avg. | |||||||||||||||||||||||
| LLM-QAT (SI8–W4) | 1B | |||||||||||||||||||||||||
| LLM-QAT (SI8–W4) | 10B | |||||||||||||||||||||||||
| LLM-QAT (SI8–W4) | 20B | |||||||||||||||||||||||||